OpenStudy (anonymous):

a pebble is dropped from the to of a building into a pool at ground level. how high is the building if the splash is seen 9 seconds after the weight is dropped. what is the velocity of the weight when it hits the ground?

OpenStudy (anonymous):

Vi=0 Vf=? t=9s g=9.81m/s x=(vi)t+(1/2)at^2 Vf=vi+at

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

The g and a variable are interchangeable and remember that acceleration is meters per second squared.